Find the area of the region between the curves

y=e3x andy=e−x from x=−1 to x=1Flhd the area of the region between the curves! ve mul v=e* from 3-10 = 1 y = €3x and y=e- from x = -1 to x = 1 Area = | 21.6
Answer #1

We have to find the shooded area show = of e30 + S et de = M - = 6 - eto-let-1) 호 일 sų -0.01656 - 0-3678 El-0.38436 Tare =0.If you are satisfied with the solution, can upvote ? ?